# The inverse Kalman filter

### v3

@@ Abstract
We introduce the inverse Kalman filter (IKF), which enables exact matrix-vector multiplication between a covariance matrix from a dynamic linear model and any real-valued vector with linear computational cost. We integrate the IKF with the conjugate gradient algorithm, which substantially accelerates the computation of matrix inversion for a general form of covariance matrix, where other approximation approaches may not be directly applicable. We demonstrate the scalability and efficiency of the IKF approach through applications in nonparametric estimation of particle interaction functions, using both simulations and real cell trajectory data.
@@ 1 Introduction
where $\mathbf{F}_{t}$ and $\mathbf{G}_{t}$ are matrices of dimensions $1\times q$ and $q\times q$, respectively, $\mathbf{W}_{t}$ is a $q\times q$ covariance matrix for $t=2,\dots,N$, and the initial state vector $\bm{\theta}_{1}\sim\mathcal{MN}(\mathbf{b}_{1},\mathbf{W}_{1})$ with $\mathbf{b}_{1}=\mathbf{0}$ assumed herein. The dependent structure of DLMs is illustrated in Fig. 1(a), where each observation $y_{t}$ is associated with a single latent state $\bm{\theta}_{t}$.
DLMs are a flexible class of models that include many widely used processes, such as autoregressive and moving average processes [26, 27]. Some Gaussian processes (GPs) with commonly used kernel functions, including the Matérn covariance with a half-integer roughness parameter [12], can also be represented as DLMs [38, 13], with closed-form expressions for $\mathbf{F}_{t}$, $\mathbf{G}_{t}$ and $\mathbf{W}_{t}$. This connection, summarized in the Supplementary Material, enables differentiable GPs within the DLM framework, allowing efficient estimation of smooth functions from noisy observations.
The Kalman filter (KF) provides a scalable approach for estimating latent state and computing likelihood in DLMs, which scales linearly with the number of observations [16], as reviewed in Appendix A. In particular, the KF enables efficient computation of $\mathbf{L}^{-1}\mathbf{u}$ for any $N$-dimensional vector $\mathbf{u}$ in $\mathcal{O}(q^{3}N)$ operations, where $\mathbf{L}$ is the Cholesky factor of the covariance matrix $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]=\mathbf{L}\mathbf{L}^{T}$, summarized in Lemma 6 in the Appendix.
@@
Computational challenges remain for high-dimensional state space models and scenarios where KF cannot be directly applied, such as the dependent structure shown in Fig. 1(b), where each observation $y_{i}$ can be associated with multiple latent states. This interaction structure, introduced in Section 3, is common in physical models, such as molecular dynamics simulation [28]. One way to overcome these challenges is to efficiently compute $\bm{\Sigma}\mathbf{u}$ for any $N$-dimensional vector $\mathbf{u}$ and utilize optimization methods such as the conjugate gradient (CG) algorithm for computing the predictive distribution [15]. This strategy has been used to approximate the maximum likelihood estimator of parameters in GPs [32, 25]. Yet, each CG iteration requires matrix-vector multiplication, which involves $\mathcal{O}(N^{2})$ operations and storage, making it prohibitive for large $N$.
To address this issue, we introduce the inverse Kalman filter (IKF), which computes $\mathbf{L}\mathbf{u}$ and $\mathbf{L}^{T}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$ operations without approximations, where $\mathbf{L}$ is the Cholesky factor of any $N\times N$ DLM-induced covariance $\bm{\Sigma}$ and $\mathbf{u}$ is an $N$-dimensional vector. The complexity is significantly smaller than $\mathcal{O}(N^{2})$ in direct computation. The latent states dimension $q$ is not larger than 3 in our applications, and it includes many commonly used models such as twice differentiable GPs with a Matérn covariance in (15), often used as a default choice in GP surrogate models.
The IKF algorithm can be extended to accelerate matrix multiplication and inversion, a key computational bottleneck in many problems. We integrate the IKF into a CG algorithm to scalably compute the predictive distribution of observations with a general covariance structure:
@@
where $\bm{\Sigma}_{j}$ is a DLM-induced covariance matrix, $\mathbf{A}_{j}$ is sparse, and $\bm{\Lambda}$ is a diagonal matrix. This structure appears in nonparametric estimation of particle interactions, which will be introduced in Section 3. Both real and simulated studies involve numerous particles over a long time period, where conventional approximation methods [34, 19, 3, 7, 10] are not applicable. This covariance structure also appears in other applications, including varying coefficient models [14] and estimating incomplete lattices of correlated data [31]. We apply our approach to the latter in Section S5 and provide numerical comparisons with various approximation methods in Section S9 of the Supplementary Material.
@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in Eqs. (1)-(2). This algorithm is applicable to any DLM specified in Eqs. (1)-(2). Denote $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We provide fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th element of the vector of $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
The algorithm in Lemma 1 is unconventional and nontrivial to derive. In contrast, Lemma 2 has direct connection to the Kalman filter. Specifically, the one-step-ahead predictive distribution in step (ii) of the Kalman filter, given in Appendix A, enables computing $\mathbf{L}^{-1}\mathbf{x}$ for any vector $\mathbf{x}$ (Lemma 6). Lemma 2 essentially reverses this operation, computing $\mathbf{L}\mathbf{\tilde{x}}$ for any vector $\mathbf{\tilde{x}}$, leading to the term Inverse Kalman filter (IKF).
The IKF algorithm, outlined in Algorithm 1, sequentially applies Lemmas 1 and 2 to reduce the computational cost and storage for $\bm{\Sigma}{\mathbf{u}}$ from $\mathcal{O}(N^{2})$ to $\mathcal{O}(q^{3}N)$ without approximations. This approach applies to all DLMs, including GPs that admit equivalent DLM representations, such as GPs with Matérn kernels where roughness parameters $0.5$ and $2.5$ correspond to $q=1$ and $q=3$, respectively. Details are provided in the Supplementary Material.
@@
2:Use Kalman filter from Lemma 5 to compute $\mathbf{B}_{t}$, $Q_{t}$, and $\mathbf{K}_{t}$ for $t=1,\dots,N$.
@@
Figure 2: Panels (a) and (b) compare the IKF (blue dots) with direct computation (red triangles) for computing $\bm{\Sigma}\mathbf{u}$ in the original time scale and logarithmic scale (base 10), respectively. Panel (c) shows the predictive error comparison between non-robust IKF (light blue squares) and robust IKF (blue dots).
While Algorithm 1 can be applied to DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in Eq. (1). This is because when $V_{t}=0$, $Q_{t}$ in Eq. (27), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in Eq. (28) due to unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the results of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
We compare the IKF approach with the direct matrix-vector multiplication of $\bm{\Sigma}\mathbf{u}$ in noise-free scenarios in Fig. 2. The experiments utilize $z_{t}=z(d_{t})$ with $d$ uniformly sampled from $[0,1]$ and the Matérn covariance with unit variance, roughness parameter $2.5$, and range parameter $\gamma=0.1$ [12]. Panels (a) and (b) show that IKF significantly reduces computational costs compared to direct computation, with a linear relationship between computation time and the number of observations. IKF achieves covariance matrix-vector multiplication for a $10^{6}$-dimensional output vector in about 2 seconds on a desktop, making it highly efficient for iterative algorithms like the CG algorithm [15] and the randomized log-determinant approximation [29]. Fig. 2(c) compares the maximum absolute error between robust IKF with $V=0.1$ and non-robust IKF with $V=0$. The non-robust IKF exhibits large numerical errors due to instability when $Q_{t}$ approaches zero, while robust IKF remains stable by ensuring $Q_{t}$ is no smaller than $V$, even with near-singular covariance matrices $\bm{\Sigma}$.
@@ 3 Nonparametric estimation of particle interaction functions
The IKF algorithm is motivated by the problem of estimating interactions between particles, which is crucial for understanding complex behaviors of molecules and active matter, such as migrating cells and flocking birds. Minimal physical models, such as the Vicsek model [35] and their extensions [1, 9], provide a framework for explaining the collective motion of active matter.
Consider a physical model encompassing multiple types of latent interactions. Let $\mathbf{y}_{i}$ be a $D_{y}$-dimensional output vector representing the $i$th particle, which is influenced by $J$ distinct interaction types. For the $j$th type of interaction, the $i$th particle interacts with a subset $p_{i,j}$ of particles rather than all particles, typically those within a radius $r_{j}$. This relationship is expressed as a latent factor model:
@@
where $\bm{\epsilon}_{i}\sim\mathcal{MN}(0,\sigma^{2}_{0}I_{D_{y}})$ denotes a Gaussian noise vector. The term $\mathbf{a}_{i,j,k}$ is a $D_{y}$-dimensional known factor loading that links the $i$th output to the $k$th neighbor in the $j$th interaction, with the unknown interaction function $z_{j}(\cdot)$ evaluated at a scalar-valued input $d_{i,j,k}$, such as the distance between particles $i$ and $k$, for $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, and $j=1,\dots,J$. For a dataset of $n_{p}$ particles over $n_{\tau}$ time points, the total number of observations is $\tilde{N}=nD_{y}=n_{p}n_{\tau}D_{y}$.
An illustrative example of this framework is the Vicsek model, a seminal approach for studying collective motion [35] (see Fig. S1(a) in the Supplementary Material). In this model, the 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, has a constant magnitude $v=||\mathbf{v}_{i^{\prime}}(\tau)||$, where |||||$\cdot$||| | ⋅ | | denotes the $L_{2}$ norm, and $v_{i^{\prime},1}(\tau)$ and $v_{i^{\prime},2}(\tau)$ are the components of the velocity along two orthogonal directions for $i^{\prime}=1,\dots,n_{p}$ and $\tau=1,\dots,n_{\tau}$. The velocity angle $\phi_{i^{\prime}}(\tau)=\tan^{-1}(v_{i^{\prime},2}(\tau)/v_{i^{\prime},1}(\tau))$ is updated as
$\phi_{i^{\prime}}(\tau)$ $=\frac{1}{p_{i^{\prime}}(\tau-1)}{\sum_{k\in ne_{i^{\prime}}(\tau-1)}\phi_{k}(\tau-1)}+\epsilon_{i^{\prime}}(\tau),$ (14)
@@
where $\epsilon_{i^{\prime}}(\tau)$ is a zero-mean Gaussian noise with variance $\sigma^{2}_{0}$. The set of neighbors $ne_{i^{\prime}}(\tau-1)$ includes particles within a radius of $r$ from particle $i^{\prime}$ at time $\tau-1$, i.e., $ne_{i^{\prime}}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r\}$, where $\mathbf{s}_{i^{\prime}}(\tau-1)$ and $\mathbf{s}_{k}(\tau-1)$ are 2-dimensional position vectors of particle $i^{\prime}$ and its neighbor $k$, respectively, and $p_{i^{\prime}}(\tau-1)$ denotes the total number of neighbors of the $i^{\prime}$th particle, including itself, at time $\tau-1$.
@@
The Vicsek model in Eq. (14) is a special case of the general framework in Eq. (13) with one-dimensional output $y_{i}=\phi_{i^{\prime}}(\tau)$, i.e. $D_{y}=1$, with the index of the $i$th observation $i=(\tau-1)n_{p}+i^{\prime}$ being a function of time $\tau$ and particle $i^{\prime}$. The Viscek model contains a single interaction, i.e. $J=1$, with linear interaction function $z(d)=d$, where $d$ being the velocity angle of the neighboring particle, and the factor loading being $1/p_{i^{\prime}}(\tau-1)$.
@@
Numerous other physical and mathematical models of self-propelled particle dynamics can also be formulated as in Eq. (13) with a parametric form of a particle interaction function [1]. Nonparametric estimation of particle interactions is preferred when the physical mechanism is unknown [17, 23], but the high computational cost limits its applicability to systems with large numbers of particles over long time periods.
@@
In this work, we model each latent interaction function $z_{j}(\cdot)$ nonparametrically using a GP. By utilizing kernels with an equivalent DLM representation, we can leverage the proposed IKF algorithms to significantly expedite computation. This includes commonly used kernels such as Matérn kernel with a half-integer roughness parameter [12], often used as the default setting of the GP models for predicting nonlinear functions, as the smoothness of the process can be controlled by its roughness parameters [11].
@@
$c_{j}(d)=\left(1+\frac{{5}^{1/2}d}{\gamma_{j}}+\frac{5d^{2}}{3\gamma_{j}^{2}}\right)\exp\left(-\frac{{5}^{1/2}d}{\gamma_{j}}\right)\,.$ (15)

### v4

@@ Abstract
We introduce the inverse Kalman filter, which enables exact matrix-vector multiplication between a covariance matrix from a dynamic linear model and any real-valued vector with linear computational cost. We integrate the inverse Kalman filter with the conjugate gradient algorithm, which substantially accelerates the computation of matrix inversion for a general form of covariance matrix, where other approximation approaches may not be directly applicable. We demonstrate the scalability and efficiency of the proposed approach through applications in nonparametric estimation of particle interaction functions, using both simulations and cell trajectories from microscopy data.
@@ 1 Introduction
where $\mathbf{F}_{t}$ and $\mathbf{G}_{t}$ are matrices of dimensions $1\times q$ and $q\times q$, respectively, $\mathbf{W}_{t}$ is a $q\times q$ covariance matrix for $t=2,\dots,N$, and the initial state vector $\bm{\theta}_{1}\sim\mathcal{MN}(\mathbf{b}_{1},\mathbf{W}_{1})$ with $\mathbf{b}_{1}=\mathbf{0}$ assumed herein. The dependent structure of DLMs is illustrated in Fig. 1(a).
DLMs are a large class of models that include many widely used processes, such as autoregressive and moving average processes [27, 28]. Some Gaussian processes with commonly used covariance functions, including the Matérn covariance with a half-integer roughness parameter [13], can also be represented as DLMs [39, 14], with closed-form expressions for $\mathbf{F}_{t}$, $\mathbf{G}_{t}$ and $\mathbf{W}_{t}$. This connection, summarized in the Supplementary Material, enables differentiable Gaussian processes to be used for scalably estimating smooth latent functions from noisy observations.
The Kalman filter (KF) provides a fast approach for estimating latent states and computing likelihood functions in DLMs, which scales linearly with the number of observations [17], as reviewed in Appendix A. In particular, the KF enables efficient computation of $\mathbf{L}^{-1}\mathbf{u}$ for any $N$-dimensional vector $\mathbf{u}$ in $\mathcal{O}(q^{3}N)$ operations, where $\mathbf{L}$ is the Cholesky factor of the covariance matrix $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]=\mathbf{L}\mathbf{L}^{T}$, summarized in Lemma A2 in the Appendix.
@@
Computational challenges remain for high-dimensional state space models and scenarios where the KF cannot be directly applied, such as the dependent structure shown in Fig. 1(b), where each observation $y_{i}$ can be associated with multiple latent states. This interaction structure, introduced in Section 3, is common in physical models, including molecular dynamics simulation [29]. A way to overcome these challenges is to efficiently compute $\bm{\Sigma}\mathbf{u}$ for any $N$-dimensional vector $\mathbf{u}$ and utilize optimization methods such as conjugate gradient algorithms for computing predictive distributions [16]. This strategy has been used to approximate the maximum likelihood estimator of parameters in Gaussian processes [33, 26]. Yet, each conjugate gradient iteration requires matrix-vector multiplication, which involves $\mathcal{O}(N^{2})$ operations and storage, making it prohibitive for large $N$.
To address this issue, we introduce the inverse Kalman filter (IKF), which computes $\mathbf{L}\mathbf{u}$ and $\mathbf{L}^{T}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$ operations without approximation, where $\mathbf{L}$ is the Cholesky factor of any $N\times N$ DLM-induced covariance $\bm{\Sigma}$ and $\mathbf{u}$ is an $N$-dimensional vector. The complexity is significantly smaller than $\mathcal{O}(N^{2})$ in direct computation. The dimension of the latent states $q$ does not exceed 3 in our applications, including commonly used models such as twice differentiable Gaussian processes with Matérn covariance in (15), often used as default in Gaussian process models.
The IKF algorithm can be extended to accelerate matrix multiplication and inversion, a key computational bottleneck in practice. We integrate the IKF into the conjugate gradient algorithm to scalably compute predictive distributions of observations with a general covariance structure:
@@
where $\bm{\Sigma}_{j}$ is a DLM-induced covariance matrix, $\mathbf{A}_{j}$ is sparse, and $\bm{\Lambda}$ is a diagonal matrix. This structure appears in nonparametric estimation of particle interactions, which will be introduced in Section 3. Both real and simulated studies involve numerous particles over a long time period, where conventional approximation methods [35, 20, 23, 3, 7, 10] are not applicable. This covariance structure also appears in other applications, including varying coefficient models [15] and estimating incomplete lattices of correlated data [32]. Though we focus on nonparametric estimation of particle interaction functions in this work, we also apply our approach to the latter in Section S5 and provide numerical comparisons with various approximation methods in Section S9 of the Supplementary Material.
@@ 2 Inverse Kalman filter
In this section, we introduce an exact algorithm for computing $\bm{\Sigma}\mathbf{u}$ with $\mathcal{O}(q^{3}N)$, where $\mathbf{u}$ is an $N$-dimensional real-valued vector and $\bm{\Sigma}=\mbox{cov}[\mathbf{y}_{1:N}]$ is an $N\times N$ covariance matrix of observations induced by a DLM with $q$ latent states in Eqs. (1)-(2). This algorithm is applicable to any DLM specified in Eqs. (1)-(2). Denote $\bm{\Sigma}=\mathbf{L}\mathbf{L}^{T}$, where the Cholesky factor $\mathbf{L}$ does not need to be explicitly computed in our approach. We develop fast algorithms to compute $\mathbf{\tilde{x}}=\mathbf{L}^{T}\mathbf{u}$ and $\mathbf{x}=\mathbf{L}\mathbf{\tilde{x}}$ by Lemma 1 and Lemma 2, respectively, each with $\mathcal{O}(q^{3}N)$ operations. Detailed proofs of these lemmas are available in the Supplementary Material. In the following, $u_{t}$, $x_{t}$, $\tilde{x}_{t}$ denote the $t$th element of the vector of $\mathbf{u}$, $\mathbf{x}$, and $\mathbf{\tilde{x}}$, respectively, and $L_{t^{\prime},t}$ denotes the $(t^{\prime},t)$th element of $\mathbf{L}$ for $t$ and $t^{\prime}=1,2,\dots,N$.
@@
The algorithm in Lemma 1 is unconventional and nontrivial to derive. In contrast, Lemma 2 has a direct connection to the Kalman filter. Specifically, the one-step-ahead predictive distribution in step (ii) of the Kalman filter, given in Appendix A, enables computing $\mathbf{L}^{-1}\mathbf{x}$ for any vector $\mathbf{x}$ (Lemma A2). Lemma 2 essentially reverses this operation, computing $\mathbf{L}\mathbf{\tilde{x}}$ for any vector $\mathbf{\tilde{x}}$, leading to the term Inverse Kalman filter (IKF).
The IKF algorithm, outlined in Algorithm 1, sequentially applies Lemmas 1 and 2 to reduce the computational cost and storage for $\bm{\Sigma}{\mathbf{u}}$ from $\mathcal{O}(N^{2})$ to $\mathcal{O}(q^{3}N)$ without approximation. The IKF applies to all DLMs, including Gaussian processes that admit equivalent DLM representation, such as Gaussian processes with Matérn kernels where roughness parameters $0.5$ and $2.5$ correspond to $q=1$ and $q=3$, respectively. Details are provided in the Supplementary Material.
@@
2:Use Kalman filter from Lemma A1 to compute $\mathbf{B}_{t}$, $Q_{t}$, and $\mathbf{K}_{t}$ for $t=1,\dots,N$.
@@
Figure 2: Panels (a) and (b) compare the IKF (blue dots) with direct computation (red triangles) for computing $\bm{\Sigma}\mathbf{u}$ in the original time scale and logarithmic scale (base 10), respectively. Panel (c) shows the predictive error comparison between the non-robust IKF (light blue squares) and the robust IKF (blue dots).
While Algorithm 1 can be applied to DLM regardless of the noise variance, we do not recommend directly computing $\bm{\Sigma}\mathbf{u}$ in scenarios when the noise variance $V_{t}$ is zero or close to zero in Eq. (1). This is because when $V_{t}=0$, $Q_{t}$ in Eq. (27), the variance of the one-step-ahead predictive distribution, could be extremely small, which leads to large numerical errors in Eq. (28) due to unstable inversion of $Q_{t}$. To ensure robustness, we suggest computing $\mathbf{x}=(\bm{\Sigma}+V\mathbf{I}_{N}){\mathbf{u}}$ with a positive scalar $V$ and outputting $\mathbf{x}-V{\mathbf{u}}$ as the result of $\bm{\Sigma}{\mathbf{u}}$. Here, $\bm{\Sigma}+V\mathbf{I}_{N}$ can be interpreted as the covariance matrix of a DLM with noise variance $V_{t}=V$. Essentially, we introduce artificial noise to ensure that $Q_{t}$ computed in KF is at least $V$ for numerical stability. The result remains exact and independent of the choice of $V$.
We compare the IKF approach with the direct matrix-vector multiplication of $\bm{\Sigma}\mathbf{u}$ in noise-free scenarios in Fig. 2. The experiments utilize $z_{t}=z(d_{t})$ with $d$ uniformly sampled from $[0,1]$ and the Matérn covariance with unit variance, roughness parameter $2.5$, and range parameter $\gamma=0.1$ [13]. Panels (a) and (b) show that the IKF significantly reduces computational costs compared to direct computation. The IKF achieves covariance matrix-vector multiplication for an output vector of $10^{6}$ dimensions in about 2 seconds on a desktop, making it highly scalable for iterative algorithms like the conjugate gradient algorithm [16] and the randomized log-determinant approximation [30]. Fig. 2(c) compares the maximum absolute error between the robust IKF with $V=0.1$ and the non-robust IKF with $V=0$. The non-robust IKF exhibits large numerical errors due to instability when $Q_{t}$ approaches zero, while the robust IKF remains stable by ensuring $Q_{t}$ is no smaller than $V$, even with near-singular covariance matrices $\bm{\Sigma}$.
@@ 3 Nonparametric estimation of particle interaction functions
The IKF algorithm is motivated by the problem of estimating interactions between particles, which is crucial to understanding complex behaviors of molecules and active matter, such as migrating cells and flocking birds. Minimal physical models, such as the Vicsek model [36] and their extensions [1, 9], provide a framework for explaining the collective motion of active matter.
Consider a physical model encompassing multiple types of latent interactions. Let $\mathbf{y}_{i}$ be a $D_{y}$-dimensional output vector corresponding to particle $i^{\prime}$ at time $\tau$, which is influenced by $J$ distinct interaction types. For the $j$th type of interaction, the $i$th observational vector interacts with a subset $p_{i,j}$ of particles rather than with all particles, typically those within a radius $r_{j}$. This relationship is expressed as a latent factor model:
@@
where $\bm{\epsilon}_{i}\sim\mathcal{MN}(0,\sigma^{2}_{0}I_{D_{y}})$ denotes a Gaussian noise vector. The term $\mathbf{a}_{i,j,k}$ is a $D_{y}$-dimensional known factor loading that links the $i$th output to the $k$th neighbor in the $j$th interaction, with the unknown interaction function $z_{j}(\cdot)$ evaluated at a scalar-valued input $d_{i,j,k}$, such as the distance between particles $i^{\prime}$ and $k$, for $k=1,\dots,p_{i,j}$, $i=1,\dots,n$, and $j=1,\dots,J$. For a dataset of $n_{p}$ particles over $n_{\tau}$ time points, the total number of observations is $\tilde{N}=nD_{y}=n_{p}n_{\tau}D_{y}$.
An illustrative example of this framework is an unnormalized Vicsek model. Unlike the original Vicsek model [36], which enforces a constant speed for all particles, we allow velocity magnitudes to vary over time. Time-varying velocities are more sensible in many real-world scenarios, including cellular movements studied in Section 4.4. When the system approaches a high degree of confluence, the average velocity of the cells typically decreases because of cell proliferation. The 2-dimensional velocity of the $i^{\prime}$th particle at time $\tau$, $\mathbf{v}_{i^{\prime}}(\tau)=(v_{i^{\prime},1}(\tau),v_{i^{\prime},2}(\tau))^{T}$, is updated by averaging the velocity of its neighbors from the previous time step. Specifically, for each direction $l=1,2$,
$v_{i^{\prime},l}(\tau)=\frac{1}{p_{i^{\prime}}(\tau-1)}\sum_{k\in ne_{i^{\prime}}(\tau-1)}v_{k,l}(\tau-1)+\epsilon_{i^{\prime},l}(\tau),$ (14)
@@
where $\epsilon_{i^{\prime}}(\tau)$ is a zero-mean Gaussian noise with variance $\sigma^{2}_{0}$. The set of neighbors $ne_{i^{\prime}}(\tau-1)$ includes particles within a radius of $r$ from particle $i^{\prime}$ at time $\tau-1$, i.e., $ne_{i^{\prime}}(\tau-1)=\{k:||\mathbf{s}_{i^{\prime}}(\tau-1)-\mathbf{s}_{k}(\tau-1)||<r\}$, where $\mathbf{s}_{i^{\prime}}(\tau-1)$ and $\mathbf{s}_{k}(\tau-1)$ are 2-dimensional position vectors of particle $i^{\prime}$ and its neighbor $k$, respectively, and $p_{i^{\prime}}(\tau-1)$ denotes the total number of neighbors of the $i^{\prime}$th particle, including itself, at time $\tau-1$. The illustration of this model is shown in Fig. S1(b) in the Supplementary Material.
@@
The unnormalized Vicsek model in Eq. (14) is a special case of the general framework in Eq. (13) with scalar output $\mathbf{y}_{i}=v_{i^{\prime},l}(\tau)$, i.e. $D_{y}=1$, and the index of the $i$th observation $i=(\tau-1)n_{p}+i^{\prime}$ is a function of time $\tau$ and particle $i^{\prime}$. This model contains a single interaction, i.e. $J=1$, with a linear interaction function $z(d)=d$, where $d$ is the velocity component of the neighboring particle, and the factor loading being $1/p_{i^{\prime}}(\tau-1)$.
@@
Numerous other physical models of self-propelled particles can also be formulated as in Eq. (13) with a parametric form of a particle interaction function [1]. Nonparametric estimation of particle interactions is preferred when the physical mechanism is unknown [18, 24], but the high computational cost limits its applicability to systems with large numbers of particles over long time periods.
@@
In this work, we model each latent interaction function $z_{j}(\cdot)$ nonparametrically using a Gaussian process. By utilizing an equivalent DLM representation, we can leverage the proposed IKF algorithm to significantly expedite computation. This includes commonly used kernels such as Matérn kernel with a half-integer roughness parameter [13], often used as the default setting for Gaussian process models to predict nonlinear functions, as the smoothness of the process can be controlled by its roughness parameters [12].
@@
$c_{j}(d)=\left(1+\frac{{5}^{1/2}d}{\gamma_{j}}+\frac{5d^{2}}{3\gamma_{j}^{2}}\right)\exp\left(-\frac{{5}^{1/2}d}{\gamma_{j}}\right)\,,$ (15)